OpenStudy (anonymous):

Solve the following systems using substitution. -3x+y=-2 -x+y=6

OpenStudy (ajprincess):

-3x+y=-2 -x+y=6 y=6+x -3x+6+x=-2 -3x+x=-2-6 can u solve for x nw?

OpenStudy (anonymous):

x=2?

OpenStudy (ajprincess):

nope x=4

OpenStudy (ajprincess):

nw put this value in y=6+x and find the value of y.

OpenStudy (anonymous):

y=10?

OpenStudy (ajprincess):

yes.

OpenStudy (anonymous):

and that's all?

OpenStudy (ajprincess):

yes..